find the missing numbers. the cost is $350.00 and the mark up rate is 70%. What is the selling price and mark up amount?
Answer by stanbon(75887)

find the missing numbers. the cost is $350.00 and the mark up rate is 70%. What is the selling price and mark up amount?
---
mark up = 0.7*350 = $245
-------------------------------
selling price = 350 + 245 = $595
===================================
Cheers,
Stan H.